OpenStudy (anonymous):

please please help me! |x-2|≥2x-1 thank you! i'm so confused! :(

OpenStudy (anonymous):

thank you for noticing my post, amistre64!

OpenStudy (amistre64):

an absolute value can never be negative; so try to determine the right side of it according to 0 or positive results to narrow the odds

OpenStudy (amistre64):

lol, thnks for posting it :)

OpenStudy (amistre64):

2x-1 >= 0 is a good start

OpenStudy (anonymous):

what do i do after x≥1/2 ? the answer is said to be x≤1 i'm really sorry. i'm so lost!! if the question just had an equal sign instead, i'd totally answer it but it isn't! :(

OpenStudy (amistre64):

you work it just like it had an equal sign, and then adjust it in the end really

OpenStudy (amistre64):

so we know x has to be greater than or equal to 1/2 so now solve it as tho it had an equals sign and get rid of all the values less than 1/2

OpenStudy (amistre64):

|x-2|≥2x-1 hmmm, also; the right side appears to be able to go negative since the | | parts is always greater then a negative hmmm

OpenStudy (amistre64):

-(x-2) >= 2x-1 (x-2) >= 2x-1 solve for both equations and we should be closer to an answer :)

terenzreignz (terenzreignz):

Hey, @amistre64 @dydlf might I suggest an alternative way? Let's look at a simpler case for absolute value inequalities: Suppose |x| > 5 Then for sure, any value that satisfies x > 5 would suffice, but is that the only way? It can be shown that if x < -5, the inequality |x| > 5 will still hold :) Basically, the way to solve avi's (absolute value inequalities) such as this, is, as @amistre64 said, treat it as you would an equal sign (almost, lol, but more on that on request :D ) So, in general, if you have |a| > b then.... a > b OR a < -b (remove the absolute value sign, reverse inequality, negate the other side) Our example here is just a fancier case :D

OpenStudy (amistre64):

good information terez

terenzreignz (terenzreignz):

thanks :)

OpenStudy (amistre64):

-(x-2) >= 2x-1 ; distribute the - thru -x+2 >= 2x-1 ; then solve as usual +x+1 +x+1 -------------- 3 >= 3x ; 1 >= x (x-2) >= 2x-1 ; this is just a usual result, undo the ( ) and work it x-2 >= 2x-1 -x+1 -x+1 ------------- -1 >= x how we have 2 cases to chk

OpenStudy (amistre64):

mark these results on a number line

OpenStudy (amistre64):

since we have a "step" where they differ; test a value in that area to see if it works; if it does we are good. Id use 0 as a test

terenzreignz (terenzreignz):

Don't we just take the union of the two sets? That way, since one is a subset of the other, the solution is just the superset?

OpenStudy (amistre64):

if it was an = sign, maybe; but its good to test all areas to be certain

OpenStudy (amistre64):

union means or doesnt it lol

OpenStudy (amistre64):

i like dbl chking to be sure since i can never remember the fancier ways

terenzreignz (terenzreignz):

Double checking never killed anyone, lol :D

OpenStudy (anonymous):

solve these equations +(x-2)>=2x-1; -(x-2)>=2x-1;.....